OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. The sum of the length and girth of a rectangular box to hold Harriet’s gift needs to be more than 110 inches. If the length needs to be three times the girth, what should the girth be?

OpenStudy (anonymous):

Let L = length Let G = girth L+G must be more than 110 inches L must be 3 times larger than G, or L=3G 3G + G = 110 4G = 110 G = 110/4 G = 27.5 inches L = 3G = 3 x 27.5 = 82.5 Answer: The girth must be 27.5 inches